# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ V Experimental implementation
This experimental case demonstrates the capability of MLCB to improve characterization accuracy in practical scenarios. However, it also underscores the challenges posed by time drifts and statistical uncertainties, which can reduce the benefits of enhanced learnability provided by MLCB. In this respect, we observe that the statistical uncertainties associated with $\mu_{q}^{BG}$ are significantly larger than the ones associated with $\tilde{\mu}_{q}^{BG}$. Were this not the case, the higher accuracy enabled by MLCB would have been even more impactful and apparent. The disparity in the uncertainties arises because $\tilde{\mu}_{q}^{BG}$ is derived almost directly from a small subset of CB data, whereas $\mu_{q}^{BG}$ is calculated from the directly measured $c_{4}$ value through an equivalence relation that involves multiple additional CB measurements (as detailed in Tab. 5 in Appendix B).
While this issue warrants further exploration, it is not a critical limitation for several reasons. First, in this comparison, the calculation of $\mu_{q}^{BG}$ was chosen for convenience, enabling a direct comparison with $\tilde{\mu}_{q}^{BG}$. From a learnability perspective, however, the directly measured $c_{4}$—which has a significantly smaller statistical uncertainty—is equally valid and fully exposes the advantages of MLCB. Moreover, the statistical uncertainty of $c_{4}$ could be further reduced by optimizing resource allocation, such as dedicating more experimental shots to MLCB runs. Lastly, as noted earlier, the square graph analyzed here represents one of the most challenging cases in terms of connecting MLCB outcomes with the fidelity ratio $\mu_{q}^{BG}$. For other graph types, this connection is more straightforward and introduces a much smaller increase in statistical uncertainty.
VI Improved error characterization
Having proved that MLCB can significantly reduce the number of unlearnable DOF in realistic scenarios, this raises the question to which extent this additional knowledge can improve the overall accuracy of the fitted SPL model. Note that this is not a trivial step, as it requires combining high-accuracy data, obtained from CB and MLCB, with low-accuracy ones, necessary to make up for the limited learnability of Pauli noise.
We answer this important point by numerically simulating the characterization of a $20$-qubit IQM GarnetTM QPU [22]. In particular, we consider two different dense layer structures, depicted in the two insets of Fig. 6(a). The rightmost one corresponds to the configuration used for the experiment in Sec. V, with pairs of layers defining closed square chains, whose MLCB analysis follow the scheme shown in Fig. 4(c). By contrast, in the leftmost, the relevant chains are open, with variable lengths, and can thus be analyzed according to the schemes in Fig 4(b).
Our numerical simulations consist of five steps. At first, (i) we randomly generate a realistic SPL noise model, i.e. we associate a map $\Lambda_{L}$ to each one of the four layer $L\in\{B,G,R,O\}$ based on model parameters $\vec{\lambda}^{L}$ that mimics the behavior of realistic devices, see Appendix D for more details. (ii) From each exact noise map $\Lambda_{L}$ we compute all the (products of) fidelities that are needed to fit the SPL model. In an actual experiment, those would be obtained using conventional CB, MLCB, as well as possible alternative low-accuracy methods. (iii) We artificially add uncertainties independently to all (product of) fidelities, in the form of Gaussian noise, to simulate the statistical uncertainty associated with an actual experimental characterization. For (products of) fidelities determined via conventional CB and MLCB, we consider a fixed standard deviation $\sigma$. To simulate the implementation of low-accuracy strategies that involves performing extra measurements, like the one based on unit-depth CB-like circuits, we take into account for the additional uncertainties due to SPAM noise and the lack of multiplicative precision by considering a larger standard deviation $\sigma^{\prime}\geq\sigma$. (iv) We then use these noisy fidelities to fit two SPL models for each layer: $\Lambda_{L}^{c}$, which is based on parameters $\{\vec{\lambda}^{L}_{c}\}$ fitted using solely conventional CB data and low-accuracy strategies, and $\Lambda_{L}^{m}$, whose fitted parameters $\{\vec{\lambda}^{L}_{m}\}$ also leverages the additional knowledge associated with MLCB experiments. (v) Finally, we compute the $L_{1}$ distances $\Delta_{t}$, with $t=c,m$, between the exact and the reconstructed noise models, summed over all the layers $L\in\{B,G,R,O\}$ as
$\Delta_{t}=\sum_{L}||\vec{\lambda}^{L}-\vec{\lambda}^{L}_{t}||_{1}.$ (20)
@@ VI Improved error characterization
Figure 6: (a) Reduction of the $L_{1}$ distance between the exact and characterized noise model enabled by MLCB. We consider $200$ randomly generated noise models and plot the ratio $r=\Delta_{m}/\Delta_{c}$, where the conventional strategy $c$ relies on the symmetry assumption, for each of the two layer configurations shown in the insets. (b) Similarly to the previous panel, we plot ratios $r$ associated with conventional strategies that allow to measure unlearnable quantities with accuracy $\sigma^{\prime}\geq\sigma=10^{-4}$, for the configuration shown in left inset of panel (a).
The most intricate aspect of this otherwise straightforward procedure lies in step (iv), particularly the role of fidelity products measured with MLCB during the model fitting process. Several strategies can be adopted to address this step. One straightforward method is to treat experimental data from conventional CB, MLCB, and low-accuracy methods on the same ground, by generalizing Eq. (8) to the multi-layer scenario
@@
Here, $\vec{\lambda}^{+}$ is the concatenation of the parameter vectors ${\vec{\lambda}}_{L}$, associated with each individual layer $L\in\{B,G,R,O\}$. Similarly, $\vec{\xi}^{+}$ is the concatenation of the (product of) fidelities vectors $\vec{\xi}_{L}$, which are used in the conventional approach, supplemented with the products of fidelities measured using MLCB. Eq. 21 can be then solved using non-negative least square fitting in a $4|\mathcal{K}|$-dimensional space.
@@
Alternatively, for each layer $L$, one can first identify a specific set of unlearnable fidelities $\{f^{L}_{\beta}\}$ required to fit the SPL model. For example, as discussed above, one can focus on all weight-one Pauli strings $\beta$ consisting of a single $X$ operator on the layer’s support. Since these fidelities are inherently unlearnable, their initial estimates rely on low-accuracy data. However, they can be significantly improved by imposing the high-accuracy constraints provided by MLCB, e.g., $\mu_{q}^{(L_{i}L_{j})}$. This optimization step, detailed in Appendix C, yields refined fidelity estimates $\{f^{L}_{\beta}\}$, which can then be used for fitting the SPL model layer-by-layer. In the reminder of this work, we primarily adopt this second approach, which places greater emphasis on leveraging high-accuracy MLCB data over low-accuracy estimates. However, other fitting strategies are certainly possible. A thorough exploration of such strategies represents a compelling avenue for future research, as it could unlock further improvements in model fitting accuracy and performance.
@@
The numerical results are summarized in Fig.6, where we plot the values of the ratio $r$ for different scenarios. In panel (a) we visualize the values of $r$ for $200$ randomly generated noise model, considering both the layer structures shown in the two insets. We consider a reasonable experimental uncertainty of $\sigma=10^{-4}$ [8] and we use the symmetry assumption to estimate the unlearnable individual fidelities. Regardless of the specific layer structure, we observe a $r$ consistently below $1$, showing that the possibility to measure with high accuracy extra DOF, unlocked by MLCB, leads to a better overall error characterization. In Fig. 6(b), we simulate a characterization procedure that rely on extra measurements, for example via unit-depth CB-like experiments, to estimate with low-accuracy the required unlearnable fidelities. For a fixed $\sigma=10^{-4}$, we consider different uncertainties associated with unlearnable quantities $\sigma\leq\sigma^{\prime}\leq 100\sigma$ and compute the ratio $r$ for $100$ randomly generated noise models for each value of $\sigma^{\prime}$. Provided that $\sigma^{\prime}\gtrsim 3\sigma$, which makes the distinction between high- and low-accuracy procedures meaningful and justified, we observe values of the ratio $r$ well below $1$, which saturates to a range $0.5\lesssim 0.8$ in the realistic regime $\sigma^{\prime}\gtrsim 10\sigma$. This shows again how MLCB can significantly improve the error characterization. By contrast, in the extreme and unrealistic regime $\sigma\sim\sigma^{\prime}$, MLCB can even be disadvantageous due to the amplification of the $\sigma$ error associated with the computation of ratios $\mu_{1}^{(L_{1},L_{j})}$ (discussed in the previous Section and detailed in Appendix B) that makes it more convenient to just consider the direct measurements of the unlearnable fidelities.
@@ VII Effects on error mitigation
Error characterization plays a crucial role in informing noise-aware error mitigation strategies, prompting an important question: how does MLCB and the associated improved characterization impact error mitigation performance? To address this, we conducted a series of numerical experiments analyzing the performance of PEC [13, 8, 15] on Clifford circuits. The choice of Clifford circuits is particularly advantageous because evaluating PEC performance in this context is both conceptually straightforward and computationally efficient [23]. Moreover, generic non-Clifford circuits can always be decomposed into a sum of maps associated with Clifford circuits, making the following analysis relevant, at least at a qualitative level, also in practical and general use cases. More precisely, data obtained on Clifford circuits can be indicative of worst-case scenarios performances [23].
@@
is the ratio between the exact Pauli fidelity $f^{L}_{\alpha}$, derived from $\Lambda_{L}$, and the fidelity $f_{\alpha,t}^{L}$ obtained from the characterized model $\Lambda_{L}^{t}$, with $t\in\{c,m\}$.
We consider Clifford circuits, consisting of the repeated implementations of $J$ Clifford layers $\bar{L}_{j}$, with $j\in[1,J]$. Each $\bar{L}_{i}$ is the concatenation of a layer randomly sampled from $\{B,G,R,O\}$ with a random layer consisting of 1Q Clifford gates. By preparing a $+1$ eigenstate of a Pauli operator $P_{\beta^{(0)}}$ and measuring the expectation value $O$ of its conjugation $P_{\beta^{(J)}}$ with respect to all Clifford layers, we obtain (neglecting SPAM errors and applying PEC)
@@
Here $\beta^{(i)}$ is the Pauli string obtained under conjugation by the first $i\geq 0$ Clifford layers, i.e. $P_{\beta^{(i)}}=\Phi(\prod_{j=1}^{i}U_{\bar{L}_{j}})[P_{\beta^{(0)}}]$. A perfect characterization would result in the noiseless result $O=1$ and deviation from this value are consequence of the finite accuracy of the reconstructed maps $\Lambda_{L}^{t}$.
@@ VIII Discussion
In this paper, we have introduced the MLCB protocol and showed its potential in increasing the accuracy of characterization of effective Pauli noise models, thanks to an improved learnability. Numerical simulations and a simple experiment have been performed focusing on practically relevant scenarios, featuring dense Clifford layers consisting of parallel CZ gates implemented on a QPU with square topology. In these cases, we show in detail how to set up the MLCB procedure to significantly decrease the number of unlearnable DOF while keeping the runtime overhead at a minimum. Our results showcase a scalable and resource-efficient procedure that can be readily extended to various combinations of CZ layers with minimal overhead.
While our focus has been on CZ gates and (SPL) models with weight two generators, MLCB can be generalized to other Clifford gates, higher-weight SPL models or alternative effective Pauli noise frameworks like CER [16]. These extensions, though valuable, are beyond the scope of this paper. On the other hand, MLCB is naturally compatible with generalizations of RC beyond the Pauli group with which the independent number of noise parameters can be reduced [40], as exemplified in Section II.2 for CZ gates and considering the addition of $S$ gates to the twirling set.
@@
First, we have not addressed the optimal allocation of experimental resources, such as the distribution of shots and rounds of RC, between different CB and MLCB protocols. In this regard, insights from the statistical analyses in protocols like ACES, as detailed in [12], provide valuable directions for future work. Clearly, optimizing the resource allocation could enhance the precision of characterization within a finite experimental runtime. More broadly, the study of sampling complexity associated with quantum noise characterization is an active and compelling area of research, with recent advancements detailed, e.g., in [42, 43, 41].
Second, as discussed in Section VI, efficiently fitting noise model parameters based on measured (products of) fidelities is a non-trivial task. While we propose a weighting strategy that prioritizes high-accuracy data (obtained through CB and MLCB) over lower-accuracy data from other methods, this approach, validated through numerical simulations of random noise models, is not guaranteed to be optimal. Further exploration in this area could yield more robust and efficient fitting techniques, enhancing the overall utility of MLCB.
Finally, while our implementation significantly reduces the number of unlearnable fidelities with only a minor runtime increase, we do not exclude the possibility of alternative strategies that could unlock even more DOF, potentially at the cost of increased experimental runtime. In this respect, during the finalization of this work, we became aware of theoretical advancements in [44], which propose a general framework for computing the learnability of Pauli noise models, considering also the SPAM parameters as target of learning. Our results are consistent with their findings in overlapping cases, such as the characterization of quasi-local noise on $n$-qubit systems arranged in 1D rings with two dense layers of parallel CZ gates, thus suggesting the optimality of MLCB in this scenario. In general, we believe our work complements the theoretical framework in [44] by providing practical details for characterizing general layers on 2D QPU s with square topologies, thus enabling straightforward implementation on physical devices, and by quantifying the impact of MLCB’s enhanced learnability on characterization quality and its effectiveness in error mitigation within the widely adopted SPL model.
@@ IX Acknowledgment
A.C. gratefully acknowledges S. Chen and A. Kandala for their insightful discussions during the APS March Meeting 2023, where the foundational ideas of this work have been presented. The authors acknowledge support from the German Federal Ministry of Education and Research (BMBF) under Q-Exa (grant No. 13N16062) and QSolid (grant No. 13N16161). The authors also acknowledge the entire IQM Technology team for their support in the development of this work.
Appendix A The example of a single CZ gate
@@
Let us focus on a system with $n=2$ qubits and on a layer $C$ consisting of a single CZ gate between the two qubits, a simple-yet-instructive example to discuss the limited learnability of Pauli noise. The fifteen non-trivial Pauli operators $P_{\alpha}\in\mathbb{P}_{2}$ are listed in the first row of Tab. 1 in the main text, while the second row shows their conjugation under $C$. Since there are only three orbits with cardinality one, associated with the Pauli strings $\alpha\in\{IZ,ZI,ZZ\}$, the three corresponding Pauli fidelities are the only ones that can be determined with high accuracy using standard Cycle Benchmarking (CB). See, for example, the setup in Fig. 8(a) that can measure $A_{ZZ}(f_{ZZ})^{d}$ and, thus, determine with high accuracy the individual fidelity $f_{ZZ}$. All other orbits have cardinality $2$ (which is the largest possible since, in this case, $U_{C}^{2}=I$). In several cases, however, the Pauli operators belonging to the same orbit share the same two-qubit support (for example, $\mathcal{R}^{C}_{XX}=\mathcal{R}^{C}_{YY}=\{P_{XX},P_{YY}\}$). In these cases, the orbit’s cardinality can be reduced to one by using interleaved CB and, in particular, by inserting two extra $S$ gates after each execution of the layer $C$ (and then recompiling them with the following layer of single-qubit (1Q) gates). This modifies the associated unitary as $U_{\bar{C}}=(S\otimes S)U_{C}$, providing us with the third row of Tab. 1 and, importantly, with four novel orbits with size one: $\mathcal{R}^{\bar{C}}_{\alpha}$ for $\alpha\in\{XX,XY,YX,YY\}$. An example of an interleaved CB circuit that allows to determine $f_{XX}$ is shown in Fig. 8(b). At this point, we are left with eight Pauli strings, which can be conveniently organized into two sets $\mathcal{S}_{A}=\{IX,IY,ZX,ZY\}$ and $\mathcal{S}_{B}=\{XI,YI,XZ,YZ\}$, whose associated orbits cannot be reduced to size one. This means that the Pauli fidelities corresponding to the elements of each set therefore cannot be determined individually with high accuracy. However, it is important to stress that, out of the $8$ DOF associated with these $8$ fidelities, the CB protocols still allows us to determine $6$ independent high-accuracy constraints on the products of fidelity pairs, thus leaving us with only $2$ unlearnable DOF. Indeed, focusing on set $\mathcal{S}_{A}$, one can determine the product $f_{IX}f_{ZX}$, as shown in Fig. 8(c), as well as $f_{IY}f_{ZY}$ and $f_{IX}f_{ZY}$ (the fourth product is not independent since $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$). The same applies to the other set $\mathcal{S}_{B}$. All in all, this example shows that a single CZ gate is associated with $2$ unlearnable DOF, which cannot be determined using CB protocols and which can be (non-unambiguously) identified with fidelities $f_{IX}$ and $f_{XI}$. These results mirror the findings discussed in Ref. [19] for a single CNOT gate.
@@
Figure 8: (a) Simple example of a CB circuit that allows for the high-accuracy determination of Pauli fidelity $f_{ZZ}$. For the sake of clarity, here we do not display the white boxes associated with RC and, instead, we show (in red) the Pauli operator which is measured at the end and how it propagates back (considering its conjugation under the CZ gates) all the way to the state preparation stage. (b) Similarly to the previous panel, we show how interleaved CB can determine the fidelity $f_{XX}$ with high accuracy. The extra purple diamonds show $S$ gates that are applied after every CZ (and eventually recompiled together with the non-displayed layer of 1Q gates) which map the Pauli operators $P_{YY}$ and $P_{XX}$ into each other. (c) In this case, CB protocols can only determine with high accuracy the product of fidelities $f_{XI}f_{XZ}$, which cannot be individually measured with high accuracy.
Appendix B Equivalence between products of fidelities

### v2

@@ VI Improved error characterization
Figure 6: (a) Reduction of the $L_{1}$ distance between the exact and characterized noise model enabled by MLCB , with respect to conventional CB approach and symmetry assumption. We consider $200$ randomly generated noise models and plot the ratio $r=\Delta_{m}/\Delta_{c}$ (see Eq.(19)) for each of the two layer configurations shown in the insets. (b) Similarly to the previous panel, we plot ratios $r$ associated with conventional strategies that allow measuring unlearnable quantities with accuracy $\sigma^{\prime}\geq\sigma=10^{-4}$ (e.g. via unit-depth circuits), for the configuration shown in left inset of panel (a).
The most intricate aspect of this otherwise straightforward procedure lies in step (iv), particularly the role of products of eigenvalues measured with MLCB during the model fitting process. Several strategies can be adopted to address this step. One straightforward method is to treat experimental data from conventional CB, MLCB, and low-accuracy methods on the same ground, by generalizing Eq. (6) to the multi-layer scenario
@@
Here, $\vec{\lambda}^{+}$ is the concatenation of the parameter vectors ${\vec{\lambda}}_{L}$, associated with each individual layer $L\in\{B,G,R,O\}$. Similarly, $\vec{\xi}^{+}$ is the concatenation of the (products of) eigenvalues vectors $\vec{\xi}_{L}$, which are used in the conventional approach, supplemented with the products of eigenvalues $\mathcal{F}_{\alpha}^{L_{1}L_{2}}$ measured using MLCB. Eq. 20 can be then solved using non-negative least square fitting in a $4|\mathcal{K}|$-dimensional space.
@@
Alternatively, for each layer $L$, one can first identify a specific set of unlearnable eigenvalues $\{f^{L}_{\beta}\}$ required to fit the SPL model. For example, as discussed above, one can focus on all weight-one Pauli strings $\beta$ consisting of a single $X$ operator on the layer’s support. Since these eigenvalues are inherently unlearnable, their initial estimates rely on low-accuracy data. However, they can be significantly improved by imposing the high-accuracy constraints provided by MLCB, e.g., $\mu_{q}^{(L_{i}L_{j})}$. This optimization step, detailed in Appendix E, yields refined eigenvalue estimates $\{f^{L}_{\beta}\}$, which can then be used for fitting the SPL model layer-by-layer. In the remainder of this work, we primarily adopt this second approach, which places greater emphasis on leveraging high-accuracy MLCB data over low-accuracy estimates. However, other fitting strategies are certainly possible. A thorough exploration of such strategies represents a compelling avenue for future research, as it could unlock further improvements in model fitting accuracy and performance.
@@
The numerical results are summarized in Fig.6, where we plot the values of the ratio $r$ for different scenarios. In panel (a) we visualize the values of $r$ for $200$ randomly generated noise model, considering both the layer structures shown in the two insets. We consider a reasonable experimental uncertainty of $\sigma=10^{-4}$ [46] and we use the symmetry assumption to estimate the unlearnable individual eigenvalues. Regardless of the specific layer structure, we observe a $r$ consistently below $1$, showing that the possibility to measure with high accuracy extra DOF, unlocked by MLCB, leads to a better overall error characterization. In Fig. 6(b), we simulate a characterization procedure that relies on extra measurements, for example via unit-depth CB-like experiments, to estimate with low-accuracy the required unlearnable eigenvalues. For a fixed $\sigma=10^{-4}$, we consider different uncertainties associated with unlearnable quantities $\sigma\leq\sigma^{\prime}\leq 100\sigma$ and compute the ratio $r$ for $100$ randomly generated noise models for each value of $\sigma^{\prime}$. Provided that $\sigma^{\prime}\gtrsim 3\sigma$, which makes the distinction between high- and low-accuracy procedures meaningful and justified, we observe values of the ratio $r$ well below $1$, which saturates to a range $0.5\lesssim 0.8$ in the realistic regime $\sigma^{\prime}\gtrsim 10\sigma$. This shows again how MLCB can significantly improve the error characterization. By contrast, in the extreme and unrealistic regime $\sigma\sim\sigma^{\prime}$, MLCB can even be disadvantageous due to the amplification of the $\sigma$ error associated with the computation of ratios $\mu_{1}^{(L_{1},L_{j})}$ (see App. C) that makes it more convenient to just consider the direct measurements of the unlearnable eigenvalues.
@@ VII Effects on error mitigation
Error characterization plays a crucial role in informing noise-aware error mitigation strategies, prompting an important question: how does MLCB and the associated improved characterization impact error mitigation performance? To address this, we conducted a series of numerical experiments analyzing the performance of PEC [43, 46, 32, 15] on Clifford circuits. The choice of Clifford circuits is particularly advantageous because evaluating PEC performance in this context is both conceptually straightforward and computationally efficient [22]. Moreover, generic non-Clifford circuits can always be decomposed into a sum of maps associated with Clifford circuits, making the following analysis relevant, at least at a qualitative level, also in practical and general use cases. More precisely, data obtained on Clifford circuits can be indicative of worst-case scenarios performances [22].
@@
is the ratio between the exact Pauli eigenvalue $f^{L}_{\alpha}$, derived from $\Lambda_{L}$, and the eigenvalue $f_{\alpha,t}^{L}$ obtained from the characterized model $\Lambda_{L}^{t}$, with $t\in\{c,m\}$.
We consider Clifford circuits, consisting of the repeated implementations of $J$ Clifford layers $\bar{L}_{j}$, with $j\in[1,J]$. Each $\bar{L}_{i}$ is the concatenation of a layer randomly sampled from $\{B,G,R,O\}$ with a random layer consisting of single-qubit Clifford gates. By preparing a +1+1 eigenstate of a Pauli operator $P_{\beta^{(0)}}$ and measuring the expectation value $O$ of its conjugation $P_{\beta^{(J)}}$ with respect to all Clifford layers, we obtain (neglecting SPAM errors and applying PEC)
@@
Here $\beta^{(i)}$ is the Pauli string obtained under conjugation by the first $i\geq 0$ Clifford layers, i.e. $P_{\beta^{(i)}}=\Phi(\prod_{j=1}^{i}U_{\bar{L}_{j}})[P_{\beta^{(0)}}]$. A perfect characterization would result in the noiseless result $O=1$ and deviations from this value are a consequence of the finite accuracy of the reconstructed maps $\Lambda_{L}^{t}$.
@@ VIII Discussion
In this paper, we have introduced the MLCB protocol and showed its potential in increasing the accuracy of characterization of effective Pauli noise models, thanks to improved learnability. Numerical simulations and a simple experiment have been performed focusing on practically relevant scenarios, featuring dense Clifford layers consisting of parallel CZgates implemented on a QPU with square topology. In these cases, we show in detail how to set up the MLCB procedure to significantly decrease the number of unlearnable DOF while keeping the runtime overhead at a minimum. Our results showcase a scalable and resource-efficient procedure that can be readily extended to various combinations of CZlayers with minimal overhead.
While our focus has been on CZgates and (SPL) models with weight-two generators, MLCB can be generalized to other Clifford gates, higher-weight SPL models or alternative effective Pauli noise frameworks like CER [7]. These extensions, though valuable, are beyond the scope of this paper. On the other hand, MLCB is naturally compatible with generalizations of PT beyond the Pauli group with which the independent number of noise parameters can be reduced [47], as exemplified in Section III.1 for CZgates and considering the addition of $S$ gates to the twirling set.
@@
First, we have not addressed the optimal allocation of experimental resources, such as the distribution of shots and rounds of Randomized Compiling (RC), between different CB and MLCB protocols. In this regard, insights from the statistical analyses in protocols like ACES, as detailed in [29], provide valuable directions for future work. Clearly, optimizing the resource allocation could enhance the precision of characterization within a finite experimental runtime. The implementation of PT at the hardware level, reaching the ideal limit of a single shot per randomized circuit, is also likely to significantly boost the precision of the characterization protocol with a fixed runtime[21, 23]. More broadly, the study of sampling complexity associated with quantum noise characterization is an active and compelling area of research, with recent advancements detailed, e.g., in [12, 13, 10].
Second, as discussed in Section VI, efficiently fitting noise model parameters based on measured (products of) eigenvalues is a non-trivial task. While we propose a weighting strategy that prioritizes high-accuracy data (obtained through CB and MLCB) over lower-accuracy data from other methods, this approach, validated through numerical simulations of random noise models, is not guaranteed to be optimal. Further exploration in this area could yield more robust and efficient fitting techniques, enhancing the overall utility of MLCB.
Finally, while our implementation significantly reduces the number of unlearnable eigenvalues with only a minor runtime increase, we do not exclude the possibility of alternative strategies that could unlock even more DOF, potentially at the cost of increased experimental runtime. In this respect, during the finalization of this work, we became aware of theoretical advancements in [11], which propose a general framework for computing the learnability of Pauli noise models, considering also the SPAM parameters as target of learning. Our results are consistent with their findings in overlapping cases, such as the characterization of quasi-local noise on $n$-qubit systems arranged in 1D rings with two dense layers of parallel CZgates, thus suggesting the optimality of MLCB in this scenario. In general, we believe our work complements the theoretical framework in [11] by providing practical details for characterizing general layers on 2D QPU s with square topologies, thus enabling straightforward implementation on physical devices, and by quantifying the impact of MLCB’s enhanced learnability on characterization quality and its effectiveness in error mitigation within the widely adopted SPL model.
@@ IX Acknowledgment
A.C. gratefully acknowledges S. Chen and A. Kandala for their insightful discussions during the APS March Meeting 2024, where the foundational ideas of this work have been presented. The authors acknowledge support from the German Federal Ministry of Education and Research (BMBF) under Q-Exa (grant No. 13N16062) and QSolid (grant No. 13N16161). The authors also acknowledge the entire IQM Technology team for their support in the development of this work.
Appendix A Pauli twirling
Here we briefly detail how the noise associated with the execution Clifford layer $C$ can be recast into a Pauli channel via twirling. Formally, the twirling of the map $\phi(U_{C})$ is performed by sandwiching $C$ in between two layers of single-qubit operators randomly sampled from the twirling set $\mathbb{T}$. Unless explicitly stated otherwise, in this work we focus on the standard Pauli Twirling (PT), with $\mathbb{T}$ consisting of all possible product combination of single-qubit Pauli operators, i.e. $\mathbb{T}=\mathbb{P}_{n}:=\{I,X,Y,Z\}^{\otimes n}$. Instead of executing $U_{C}$, we first implement a twirling layer $T\in\mathbb{T}$, followed by $U_{C}$ and then by the conjugated layer $T_{C}:=U_{C}TU_{C}^{\dagger}\in\mathbb{T}$, which also belongs to $\mathbb{T}$ since the Pauli group is invariant under Clifford conjugation. If we assume that the noise map $\Omega$ associated with the execution of a layer $T$ of single-qubit gates is gate-independent, i.e. $\nu(T):=\Omega\circ\phi(T)$ $\forall T$, one can show that uniformly sampling $T$ from the twirling set $\mathbb{T}$ and subsequently averaging the results, denoted as $\langle\cdot\rangle_{T\in\mathbb{T}}$, leads to the following effective evolution
where
is an effective Pauli channel associated with the execution of the Pauli-dressed Clifford layer $C$. The channel $\Lambda_{C}$ corresponds to the noise of the original Clifford layer $C$, which contributes the majority of the noise, together with the noise of the twirling layer $T$.
Appendix B The example of a single CZgate
@@
Let us provide more detail about the simple-yet-instructive example of a system with $n=2$ qubits and on a layer $C$ consisting of a single CZgate between the two qubits. The fifteen non-trivial Pauli operators $P_{\alpha}\in\mathbb{P}_{2}$ are listed in the first row of Tab. 1 in the main text, while the second row shows their conjugation under $C$. Since there are only three orbits with cardinality one, associated with the Pauli strings $\alpha\in\{IZ,ZI,ZZ\}$, the three corresponding Pauli eigenvalues are the only ones that can be determined with high accuracy using standard Cycle Benchmarking (CB) directly on the layer $C$. See, for example, the setup in Fig. 8(a) that can measure $A_{ZZ}(f_{ZZ})^{d}$ and, thus, determine with high accuracy the individual eigenvalue $f_{ZZ}$. All other orbits have cardinality $2$ (which is the largest possible since, in this case, $U_{C}^{2}=I$). In several cases, however, the Pauli operators belonging to the same orbit share the same two-qubit support (for example, $\mathcal{R}^{C}_{XX}=\mathcal{R}^{C}_{YY}=\{P_{XX},P_{YY}\}$). In these cases, the orbit’s cardinality can be reduced to one by inserting two extra $S$ gates after each execution of the layer $C$ (and then recompiling them with the following layer of single-qubit gates). This modifies the associated unitary as $U_{\bar{C}}=(S\otimes S)U_{C}$, providing us with the third row of Tab. 1 and, importantly, with four novel orbits with size one: $\mathcal{R}^{\bar{C}}_{\alpha}$ for $\alpha\in\{XX,XY,YX,YY\}$. An example of an interleaved CB circuit that allows determining $f_{XX}$ is shown in Fig. 8(b). At this point, we are left with eight Pauli strings, which can be conveniently organized into two sets $\mathcal{S}_{A}=\{IX,IY,ZX,ZY\}$ and $\mathcal{S}_{B}=\{XI,YI,XZ,YZ\}$, whose associated orbits cannot be reduced to size one. This means that the Pauli eigenvalues corresponding to the elements of each set therefore cannot be determined individually with high accuracy. However, it is important to stress that, out of the $8$ DOF associated with these $8$ eigenvalues, the CB protocols still allows us to determine $6$ independent high-accuracy constraints on the products of eigenvalue pairs, thus leaving us with only $2$ unlearnable DOF. Indeed, focusing on set $\mathcal{S}_{A}$, one can determine the product $f_{IX}f_{ZX}$, as shown in Fig. 8(c), as well as $f_{IY}f_{ZY}$ and $f_{IX}f_{ZY}$ (the fourth product is not independent since $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$). The same applies to the other set $\mathcal{S}_{B}$. All in all, this example shows that a single CZgate is associated with $2$ unlearnable DOF, which cannot be determined using CB protocols and which can be (non-unambiguously) identified with eigenvalues $f_{IX}$ and $f_{XI}$. These results mirror the findings discussed in Ref. [9] for a single CNOT gate.
@@
Figure 8: (a) Simple example of a CB circuit that allows for the high-accuracy determination of Pauli eigenvalue $f_{ZZ}$. For the sake of clarity, here we do not display the white boxes associated with RC and, instead, we show (in red) the Pauli operator which is measured at the end and how it propagates back (considering its conjugation under the CZgates) all the way to the state preparation stage. (b) Similarly to the previous panel, we show how interleaved CB can determine the eigenvalue $f_{XX}$ with high accuracy. The extra purple diamonds show $S$ gates that are applied after every CZ(and eventually recompiled together with the non-displayed layer of single-qubit gates) which map the Pauli operators $P_{YY}$ and $P_{XX}$ into each other. (c) In this case, CB protocols can only determine with high accuracy the product of eigenvalues $f_{XI}f_{XZ}$, which cannot be individually measured with high accuracy.
Appendix C Equivalence between products of eigenvalues